Prove that the points M(-2,-2), N(2,1), and L(6,4) are on a straight line using the midpoint equation.
Answer by Boreal(15235)

midpoint is (x1+x2)/2, (y1+y2)/2
(-2+6)/2=2
(-2+4)/2)=1
(2, 1) is midpoint, which is N. The shortest distance is through the given midpoint, so a straight line.